OpenStudy (anonymous):

the perimeter of a rectagular calculator is 36cm the length is cm greater than the width. find the width and the length?

OpenStudy (anonymous):

how many cm greater than the width is the length?

OpenStudy (anonymous):

the lenght is 4cm perimeter is 36 cm

OpenStudy (anonymous):

4cm greater than the width srry guys how about to write dwn the problem

OpenStudy (anonymous):

Sorry, I lost my explanation. One sec.

OpenStudy (anonymous):

Let : x = length y = width Think: perimeter = sum of the exterior sides of a particular polygon. Therefore, the perimeter of a rectangle = two lengths + two widths 2x + 2y = 36 ----- 1st equation We can make a second by looking at the 2nd half of the question. If the length is 4 cm greater than the width, the width is equal to the length - 4 cm. y = x - 4 ---- second equation

OpenStudy (anonymous):

i just get confuse on how to write dwn the problem thanxs

OpenStudy (anonymous):

Then we substitute one equation into the other to solve for one of the dimensions. For example, 2x + 2(x-4) = 36 4x -8 = 36 4x = 28 x = 7 cm Sub. x = 7 into either of the equations. y = 7 - 4 y = 3 cm

OpenStudy (anonymous):

thanks so much!! i think i need to practice more on this type of explanations

OpenStudy (anonymous):

No problem, word problems are often difficult for most students!

OpenStudy (anonymous):

Gail trip cost 1.50 per trip(one way). montly pass cost $54 express as inequality the # of trips per month Gail should make if the pass is to save her money. I need to be able to write dwn and solve it.. help thanks.